OpenStudy (anonymous):

find the lengths of both circular arcs on the unit circle connecting the point (-square root 2/2, -square root 2/2) and the point that makes an angle of 115 degrees with the positive horizontal axis.

OpenStudy (anonymous):



OpenStudy (anonymous):

one angle is 115 degrees. the other is 225 degrees

OpenStudy (anonymous):

angle between then is 110 degrees one direction and 360-110 = 250 in the other. you want the arc length and the total circumference of the circle is \[2\pi\] so solve \[\frac{110}{360}=\frac{x}{2\pi}\] for x to get the first answer \[x=\frac{2\times 110\times \pi}{360}=\frac{220\pi}{360}=\frac{11\pi}{18}\] if my arithmetic is correct

OpenStudy (anonymous):

yup, that worked out.

OpenStudy (anonymous):

ty